# The Generic Temperature Response of Large Biochemical Networks

### v1

@@ II.4 Partition sums and generating functions
We note that the number of spanning trees on a complete graph on $N+1$ vertices is given by the Cayley formula[79] as $(N+1))^{N-1}$, so it grows faster then exponential with $N$. Even though many of the rates vanish, for sufficiently complex networks one can expect the numerator and the denominator of Eq. (19) to be the sum of many such graphs. This motivates the introduction of a partition sum like quantity $Z$ summing over the spanning trees:
@@
Letting the sums in Eq. (23) and Eq. (24) run only over the non-vanishing trees, all involved rates satisfy $k_{ij}>0$, and one can parameterize them as $k_{ij}(a)=e^{aX_{i,j}}$ such that $k_{ij}(a=1)=e^{X_{i,j}}$ is the rate of interest. Applying this parametrization, one obtains for $Z$:
@@
defining $X_{\mathcal{T}}:=\sum_{k_{ij}\in E(\mathcal{T})}X_{ij}$, which is the sum over all $X_{ij}$ of a given tree. Now consider the Taylor-series of $\ln{Z}$ in $a=0$:
$\ln{Z_{\mathcal{T}}}=\sum_{n=0}^{\infty}\frac{\partial^{n}\ln{Z_{\mathcal{T}}}}{\partial^{n}a}\Big{|}_{a=0}\frac{a^{n}}{n!}$ (27)
@@
$\frac{\partial^{2}\ln{Z_{\mathcal{T}}}}{\partial^{2}a}\Big{|}_{a=0}$ $=\frac{\sum_{\mathcal{T}_{[N+1]}}X_{\mathcal{T}}^{2}}{|\mathcal{T}|}-\Big{(}\frac{\sum_{\mathcal{T}_{[N+1]}}X_{\mathcal{T}}}{|\mathcal{T}|}\Big{)}^{2}$
@@
$\frac{\partial^{n}\ln{Z_{\mathcal{T}}}}{\partial^{n}a}\Big{|}_{a=0}=\kappa_{n}^{X_{\mathcal{T}}}.$ (31)
Hence,
@@
and completely analogously for $\ln{Z_{\mathcal{F}}}$:
$\ln{Z_{\mathcal{F}}(a)}=\sum_{n=0}^{\infty}\kappa_{n}^{X_{\mathcal{F}}}\frac{a^{n}}{n!}.$ (33)
For the rates of interest ($a=1$), the logarithm of Eq. (25) becomes:
@@
Note that Eq. (34) is completely general and as such could be applied e.g. to the MAPK-system shown in Fig. 1. The cumulants are taken over the spanning trees and the expression is precise if all $k_{ij}$ are known for any $N$. Also, our treatment so far is not restricted to biochemical networks. In principle, it can also be applied to other settings described by a master equation.
@@ II.5 Limit of large system size
We now are in a position to study the limit of large networks ($N\gg 1$). Because in this case not all details can be known, at the same time we switch to a statistical description for the reaction rates $k_{ij}$. To address temperature dependance, we consider biochemical networks with rates that have an Arrhenius-like temperature dependence, that is to say, $X_{ij}=-\Delta\beta E_{ij}+\ln{K_{ij}}$, where $E_{ij}$ and $\ln{K_{ij}}$ are assumed to be independent random variables. For that case, using that
$\kappa_{n}^{X_{\mathcal{T}}}$ $=\kappa_{n}^{-\Delta\beta E_{\mathcal{T}}+\sum_{k_{ij}\in E(\mathcal{T})}\ln{K_{ij}}}$
$=(-\Delta\beta)^{n}\kappa_{n}^{E_{\mathcal{T}}}+\kappa_{n}^{\sum_{k_{ij}\in E(\mathcal{T})}\ln{K_{ij}}},$ (35)
@@
Note that backward rates do not fully enter in the spanning trees (cf. examples in appendix A) so the activation energies on the trees and on the two-tree forests are different in general. The result so far requires to know the cumulants $\kappa_{n}^{E_{\mathcal{T}}}$ and $\kappa_{n}^{E_{\mathcal{F}}}$ to get the global rate constant. A priori, there may be infinitely many of them. But there are two cases in which the description simplifies:
Case 1: All $E_{ij}$ are independent and identically distributed with no difference between the energies on the trees and the reduced trees. Then $E_{\mathcal{T}}=E_{\mathcal{F}}+E_{ij}$ (seen as random variables) and Eq. (36) becomes:
$\ln{\langle\tau\rangle}=$ $-\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!}\kappa_{n}^{E_{ij}}\Delta\beta^{n}+const.$
$=$ $\ln{\langle e^{-\Delta\beta E_{ij}}\rangle}+const,$ (37)
again using the definition of the cumulant generating function.
For a developmental process, it seems rather reasonable to assume that the clear direction in the network is also reflected by the involved activation energies, so that the distributions of the forward and backward rates is different. Thus case 1 is therefore most likely not applicable to developmental systems.
Case 2: $E_{\mathcal{T}}$ and $E_{\mathcal{F}}$ can be sufficiently well approximated by normal distributions, $E_{\mathcal{T}}\sim\mathcal{N}(\langle E\rangle_{\mathcal{T}},\sigma_{\mathcal{T}}^{2})$ and $E_{\mathcal{F}}\sim\mathcal{N}(\langle E\rangle_{\mathcal{F}},\sigma_{\mathcal{F}}^{2})$. Then all but the first two cumulants vanish, resulting in the following expression:
@@
$E_{\mathcal{T}}=\sum_{k_{ij}\in E(\mathcal{T})}E_{ij}$ and $E_{\mathcal{F}}=\sum_{k_{ij}\in E(\mathcal{F})}E_{ij}$ are the sums of $N$, and $N-1$ independent random variables and, since $N\gg 1$, the central limit theorem suggests indeed that this is a valid assumption. However, the crucial point here is that the higher cumulants have to vanish fast enough compared to the differences of the first two terms, requiring sufficiently well-behaved convergence to the normal distribution. The simplest case is that the $E_{ij}$ already follow a normal distribution, which may be different for the $E_{ij}$ on the trees and the two tree-forests, their sums also follow a normal distribution and one obtains:
@@
We thus have arrived at our main result: the Arrhenius plot for large and complex biochemical networks is expected to show a quadratic dependence on the inverse temperature, if they have a bias towards a specific target state, as it must be the case for developmental systems.
II.6 Comparison to experimental data
(a) Rate until the beginning of cellularization (stage B, $R^{2}=0.9587$)
(b) Rate until the development of horizontal posterior midgut (stage E, $R^{2}=0.9909$)
(c) Rate until full germ band retraction (stage H, $R^{2}=0.9952$)
(d) Rate until the first breath (stage K, $R^{2}=0.9868$).
@@ II.6 Comparison to experimental data
Figure 5: Arrhenius plot (logarithmic rates against inverse temperature) for the rates to the different stages (B,E,H,K) in the development of Drosophila melanogaster embryos, starting at the 14th nuclear division adapted from [26] with quadratic fit to data. The labeling of the stages follows the convention from the original authors.
(a) Male specimen ($R^{2}$=0,9950)
(b) Female specimen ($R^{2}$=0,9950)
@@
Figure 6: Logarithmic developmental rate to the pupal stage of Drosophila melanogaster versus inverse temperature adapted from [24] with quadratic fit.
(a) Embryonic development rates ($R^{2}$=0,9976)
(b) Larval development rates for male specimen ($R^{2}$=0,9838)
(c) Larval development rates for female specimen ($R^{2}$=0,9808)
(d) Pupal development rates for male specimen ($R^{2}$=0,9990)
(e) Pupal development rates for female specimen ($R^{2}$=0,9981)
@@
Figure 7: Logarithmic developmental rate for the embryonic, pupal and larval stage of Drosophila melanogaster [25] with quadratic fit.
@@
Indeed our theoretical prediction is in striking agreement with experimental data for developmental systems. In a work published in 2021, Crapse et al.[26] conducted a study on the development rates of Drosophila embryos at different temperatures. To that end, the embryonic phase of Drosophila was subdivided into 12 stages and the rates to reach each stage were measured under different temperature conditions. The 14th nuclear division in the zygote was chosen as the starting point (stage A) and the first breath was chosen as the end point (stage K). The cumulative rates starting with the 14th nuclear division to 4 different stages (B,E,H,K in the notation of the original authors) are shown in Fig. 5 as an example, together with a quadratic fit performed in Python via the numpy.poly1d command. The data were taken from the supplementary material of the original publication, which tabulated the measured times for the entire sample and all stages. The fitting parameters are listed in appendix C. Note that these do not provide insights into the microscopic structure of the network without further assumptions since the coefficients of the quadratic functions contain the differences of the first two moments along the trees and forests as well as terms containing $\frac{1}{k_{B}T_{0}}$.
@@
For the beginning of stage B (begin of cellularization), no robust statement can be made about the goodness of the quadratic fit. Since this is the first stage following the starting point, it may well be that the assumption of a sufficiently large underlying biochemical network is simply not valid. However, as more steps accumulate, such as in the rates to stages E, H, and K, the quadratic shape of the data points becomes more apparent. A quadratic fit of the data was also suggested by Crapse et al.[26], although without providing an explanation on fundamental grounds, as done here.
@@
Measurements of the temperature dependence of developmental rates of Drosophila were already performed by Bliss[24] in 1926 for the prepupal development and by Powsner[25] in 1935 for the embryonic, pupal, and larval stages. Their results are shown in Fig. 6 and 7, respectively, also with a quadratic fit to their data points. Both articles list their data in a table and show them in a diagram, albeit together with piecewise linear functions as fit instead of a global quadratic function. One sees that in all cases, the quadratic fits predicted by our theory are in very good agreement with the experimental data.

### v2

@@ II.4 Partition sums and generating functions
We note that the number of spanning trees on a complete graph on $N+1$ vertices is given by the Cayley formula [82] as $(N+1)^{N-1}$, so it grows faster than exponentially with $N$. Even though many of the rates vanish, for sufficiently complex networks one can expect the numerator and the denominator of Eq. (19) to be the sum of many such graphs. This motivates the introduction of a partition sum-like quantity $Z$ summing over the spanning trees:
@@
Letting the sums in Eq. (23) and Eq. (24) run only over the non-vanishing trees, all involved rates satisfy $k_{ij}>0$, and one can parameterize them as $k_{ij}(a)=e^{aX_{i,j}}$ such that $k_{ij}(a=1)=e^{X_{i,j}}$ is the rate of interest. Applying this parametrization, one obtains for $Z_{\mathcal{T}}$:
@@
defining $X_{\mathcal{T}}:=\sum_{k_{ij}\in E(\mathcal{T})}X_{ij}$, which is the sum over all $X_{ij}$ of a given tree. Now consider the Taylor-series of $\ln{Z_{\mathcal{T}}}$ around $a=0$:
$\ln{Z_{\mathcal{T}}}(a)=\sum_{n=0}^{\infty}\frac{\partial^{n}\ln{Z_{\mathcal{T}}}}{\partial^{n}a}\Big{|}_{a=0}\frac{a^{n}}{n!}$ (27)
@@
$\frac{\partial^{2}\ln{Z_{\mathcal{T}}}}{\partial a^{2}}\Big{|}_{a=0}$ $=\frac{\sum_{\mathcal{T}_{[N+1]}}X_{\mathcal{T}}^{2}}{|\mathcal{T}|}-\Big{(}\frac{\sum_{\mathcal{T}_{[N+1]}}X_{\mathcal{T}}}{|\mathcal{T}|}\Big{)}^{2}$
@@
$\frac{\partial^{n}\ln{Z_{\mathcal{T}}}}{\partial a^{n}}\Big{|}_{a=0}=\kappa_{n}^{X_{\mathcal{T}}}.$ (31)
Hence, one can express $\ln{Z_{\mathcal{T}}}$ as:
@@
with the convention that $\kappa_{0}^{X_{\mathcal{T}}}:=\ln{|\mathcal{T}|}$. Completely analogously for $\ln{Z_{\mathcal{F}}}$ one gets:
$\ln{Z_{\mathcal{F}}(a)}=\sum_{n=0}^{\infty}\kappa_{n}^{X_{\mathcal{F}}}\frac{a^{n}}{n!},$ (33)
with $\kappa_{0}^{X_{\mathcal{F}}}:=\ln{|\mathcal{F}|}$. For the rates of interest ($a=1$), the logarithm of Eq. (25) becomes:
@@
Note that Eq. (34) is completely general and, as such, could be applied, for example, to the MAPK-system shown in Fig. 1b. The cumulants are taken over the spanning trees and the expression is precise if all $k_{ij}$ are known for any $N$. Also, our treatment so far is not restricted to biochemical networks. In principle, it can also be applied to other settings described by a master equation.
@@ II.5 Limit of large system size
We now are in a position to study the limit of large networks ($N\gg 1$). Because in this case not all details can be known, at the same time we switch to a statistical description for the reaction rates $k_{ij}$. To describe the effect of temperature, we consider biochemical networks with rates that have an Arrhenius-like temperature dependence, that is to say, $X_{ij}=-\Delta\beta E_{ij}+\ln{k^{0}_{ij}}$, where $E_{ij}$ and $\ln{k^{0}_{ij}}$ are assumed to be independent random variables. For that case, using that
$\kappa_{n}^{X_{\mathcal{T}}}$ $=\kappa_{n}^{-\Delta\beta E_{\mathcal{T}}+\sum_{k_{ij}\in E(\mathcal{T})}\ln{k^{0}_{ij}}}$
$=(-\Delta\beta)^{n}\kappa_{n}^{E_{\mathcal{T}}}+\kappa_{n}^{\sum_{k_{ij}\in E(\mathcal{T})}\ln{k^{0}_{ij}}},$ (35)
@@
Note that backward rates do not fully enter in the spanning trees (cf. examples in Appendix A) so the activation energies on the trees and on the two-tree forests are different in general. The result so far requires to know the cumulants $\kappa_{n}^{E_{\mathcal{T}}}$ and $\kappa_{n}^{E_{\mathcal{F}}}$ to get the global rate constant.
If now $E_{\mathcal{T}}$ and $E_{\mathcal{F}}$ can be sufficiently well approximated by normal distributions, $E_{\mathcal{T}}\sim\mathcal{N}(\langle E\rangle_{\mathcal{T}},\sigma_{\mathcal{T}}^{2})$ and $E_{\mathcal{F}}\sim\mathcal{N}(\langle E\rangle_{\mathcal{F}},\sigma_{\mathcal{F}}^{2})$, then all but the first two cumulants vanish, resulting in the following expression:
@@
$E_{\mathcal{T}}=\sum_{k_{ij}\in E(\mathcal{T})}E_{ij}$ and $E_{\mathcal{F}}=\sum_{k_{ij}\in E(\mathcal{F})}E_{ij}$ are the sums of $N$, and N1N-1 independent random variables and, since $N\gg 1$, the central limit theorem suggests indeed that this is a valid assumption as long as the higher cumulants vanish fast enough compared to the differences of the first two terms, requiring sufficiently well-behaved convergence to the normal distribution. The simplest case is that the $E_{ij}$ already follow a normal distribution, which may be different for the $E_{ij}$ on the trees and the two tree-forests, their sums also follow a normal distribution and one obtains:
@@
This constitutes our main theory result: the Arrhenius plot for large and complex biochemical networks is expected to show a quadratic dependence on the inverse temperature.
II.6 Comparison to simulated networks
To illustrate the convergence predicted by theory, random networks on $N_{v}=10$ to $N_{v}=1.000$ vertices displaying some basic characteristics of the biochemical network of a developmental process were generated (see Appendix D for details) [83]. In particular, they were designed to have a relatively large distance between the initial vertex 0 and the final vertex $N_{v}$ as well as a block-like structures with many connections and irreversible transitions between them, reflecting the checkpoints present in development. Two small networks with $N_{v}=30$ and $N_{v}=50$ are shown as examples in Fig. 5a and b, respectively.
Figure 5: Randomly generated networks with features that are typical for development of sizes (a) $N_{v}=30$ and (b) $N_{v}=50$ vertices. The initial vertices are represented as squares and the final vertices as diamonds. The networks are biased towards the final vertices and contain compact elements with many connections.
Figure 6: Histograms of the total activation energies of (a) spanning trees $E_{\mathcal{T}}$ and (b) two-tree spanning forests $E_{\mathcal{F}}$, i.e., the sum of all the activation energies along their edges, on randomly generated graphs on $N_{v}=10$ vertices.
The spanning trees and two-tree spanning forests of these base networks were then drawn at random. A standard procedure to generate random spanning trees is given by Wilson’s algorithm [84]. While removing a random edge from a spanning tree will result in a spanning forest, this procedure does not yield all spanning forests with the same probability. In order to ensure an unbiased sample, a specific algorithm was devised for the generation of two-tree spanning forests. In Appendix C, the steps needed for this algorithm are explained and its functionality is illustrated for a simple example network.
Figure 7: Histograms and Arrhenius plots for simulated random networks of increasing sizes. (a-c) $N_{v}=20$. (d-f) $N_{v}=100$. (g-i) $N_{v}=1.000$ (a,d,g) Spanning trees. (b,e,h) Spanning forests. (c,f,i) Resulting Arrhenius plots. The fitted parameters are listed in Appendix B.
The activation energies were assigned randomly to the edges. The activation energies were drawn at random from the intervals $60\frac{\text{kJ}}{\text{mol}},100\frac{\text{kJ}}{\text{mol}}]$ for the forward rates and from $20\frac{\text{kJ}}{\text{mol}},60\frac{\text{kJ}}{\text{mol}}]$ for the backward rates. These represent typical values of activation energies of biochemical reactions listed in the literature[85].
In each case, $10.000$ spanning trees and at least $10.000$ two-tree spanning forests (note that the number of generated forests is not predefined to ensure an unbiased sample, as explained in Appendix C), and the total activation energies for each of them were calculated. The total activation energies for $N_{v}=10$ are shown in Fig. 6 as histograms. One observes that there are just a few spanning trees and also a moderate number of two-tree spanning forests, so that the distribution is not reminiscent of a Gaussian yet. The other histograms, along with fits to normal distributions, are shown in Fig. 7. As expected, the histograms of the total activation energies approach a normal distribution as $N_{v}$ increases.
Then, Eq. (37) can be applied to obtain the Arrhenius plot for the global rate. The corresponding plots for networks of size $20$, $100$ and $1.000$ are shown in Fig. 7. The convergence of the total energy distributions to a Gaussian is observed as the system size $N_{v}$ increases. For the parameters chosen here, the distribution appears already well described by a Gaussian for $N_{v}\sim 100$. The second derivative, i.e., twice the quadratic term as a function of $N_{v}$ is shown in Fig. (8). One observes that as the system size increases, so does the quadratic term, in a statistical manner. This proves numerically that this kind of networks generically lead to a quadratic Arrhenius plot in the limit of large system size.
Figure 8: The absolute value of the quadratic term in the Arrhenius plot $\sigma_{F}^{2}-\sigma_{T}^{2}$ versus number of vertices $N_{v}$. For every value $N_{v}$, 100 networks of this size were generated and the distribution of the total activation energies of the trees and two tree spanning forests were computed. The error bars indicate the standard deviations.
@@ II.6 Comparison to simulated networks
Figure 9: Arrhenius plot for the rates to the different stages in the development of Drosophila melanogaster embryos, starting at the 14th nuclear division (A) adapted from [24] with quadratic fit to data. The residuals to the fit are shown below the main diagram. The labeling of the stages follows the convention from the original authors: (a) rate from A to B (beginning of cellularization), $R^{2}=0.9587$ (b) rate from A to E (development of horizontal posterior midgut), $R^{2}=0.9909$ (c) rate from A to H (full germ band retraction), $R^{2}=0.9952$ (d) rate from A to K (first breath), $R^{2}=0.9868$.
@@
Figure 10: Logarithmic developmental rate to the pupal stage of Drosophila melanogaster versus inverse temperature adapted from [25] with quadratic fit. The residuals to the fit are shown below the main diagram. (a) For male specimen, $R^{2}=0.9950$ (b) for female specimen, $R^{2}=0.9950$.
@@
Figure 11: Logarithmic developmental rate for the embryonic, pupal and larval stage of Drosophila melanogaster [26] versus inverse temperature with quadratic fit. The residuals to the fit are shown below the main diagram. (a) embyonic development, $R^{2}=0.9976$ (b) male larval development, $R^{2}=0,9838$ (c) female larval development, $R^{2}=0.9808$ (d) male pupal development, $R^{2}=0.9990$ (e) female pupal development, $R^{2}=0.9981$.
Figure 12: The absolute value of the quadratic coefficient from the quadratic fit to the data from [24] with the last cleavage as starting point to later stages in development.
II.7 Comparison to experimental data
@@
Our theoretical prediction is in striking agreement with experimental data for developmental systems. In a work published in 2021, Crapse et al. [24] conducted a study on the development rates of Drosophila embryos at different temperatures. To that end, the embryonic phase of Drosophila was subdivided into 12 stages and the rates to reach each stage were measured under different temperature conditions. The 14th nuclear division in the zygote was chosen as the starting point (stage A) and the first breath was chosen as the end point (stage K). The cumulative rates starting with the 14th nuclear division to 4 different stages (B,E,H,K in the notation of the original authors) are shown in Fig. 9 as an example, together with a quadratic fit performed in Python via the numpy.polyfit command. The data were taken from the supplementary material of the original publication, which tabulated the measured times for the entire sample and all stages. The fitting parameters are listed in Appendix B. Note that these do not provide insights into the microscopic structure of the network without further assumptions since the coefficients of the quadratic functions contain the differences of the first two moments along the trees and forests as well as terms containing $\frac{1}{k_{B}T_{0}}$. Moreover, the residuals of the quadratic fit were computed to analyze whether there are any systematic deviations that hint towards a deviation from the quadratic dependence.
@@
For the beginning of stage B (begin of cellularization), no robust statement can be made about the goodness of the quadratic fit. Since this is the first stage following the starting point, it may well be that the assumption of a sufficiently large underlying biochemical network is simply not valid. However, as more steps accumulate, such as in the rates to stages E, H, and K, the quadratic shape of the data points becomes more apparent. The residuals do not point to a systematic deviation, suggesting that higher order terms at best play a minor role. A quadratic fit of the data was also suggested by Crapse et al. [24], although without providing an explanation on fundamental grounds, as done here.
Fig. 12 shows the quadratic terms of the quadratic fit for the rates to the different stages defined by Crapse et al. Consistent with the simulation data shown in Fig. 8, the quadratic term tends to increase as more steps accumulate. However, a slight decrease is observed for the rates to the final stages. This is most likely due to the lack of data points at the extreme ends of the temperature spectrum, as most embryos did not survive long enough under these conditions. Consequently, this also results in greater uncertainties for these values.
@@
Measurements of the temperature dependence of developmental rates of Drosophila were already performed by Bliss [25] in 1926 for the prepupal development and by Powsner [26] in 1935 for the embryonic, pupal, and larval stages. Their results are shown in Fig. 10 and 11, respectively, also with a quadratic fit to their data points. Both articles list their data in a table and show them in a diagram, albeit together with piecewise linear functions as fit instead of a global quadratic function. One sees that in all cases, the quadratic fits predicted by our theory are in very good agreement with the experimental data. Also the residuals indicate only a minor contributions from terms of higher order, except for the larval development rates from the data by Powsner (Fig. 11 (b,c)), where most likely some additional effect contributes.
Moreover, one also observes that a clear quadratic shape becomes more prominent the more steps accumulate, as also anticipated from the previous theoretical considerations. Note that the maximal rate, i.e. the peak in the diagrams, is at $T_{peak}\sim 30^{\circ}\text{C}$. Expecting the mean total activation energies to be higher for the trees than for the forests, and vice-versa for the variances, Eq. (37) predicts the maximum to be attained at $\Delta\beta<0$, so $T_{peak}>T_{0}$, and one would in fact assume $T_{0}\sim 20^{\circ}\text{C}$.
II.8 Deviations from the generic quadratic response
Both our theory and computer simulations demonstrate that large networks show a generic Arrhenius plot with a quadratic form. We finally discuss deviations from this generic response. We start with small networks and demonstrate that they easily can be designed to give non-linear Arrhenius plots. We then discuss the case of linear chains, which shows that even large networks can deviate, because they contain only few spanning trees.
II.8.1 Small networks
Figure 13: Example network on five vertices with a strong deviation from an Arrhenius-like response. (a) Network (b) Spanning trees (c) spanning forests (d) Arrhenius plot for the ratio $(k_{1}:k_{2}:k_{3}:k_{4}:k_{5})=(2:2:5:0.1:2)$ at $T_{0}=300K$ and $E_{1}=20\frac{\text{kJ}}{\text{mol}}$, $E_{2}=100\frac{\text{kJ}}{\text{mol}}$, $E_{3}=30\frac{\text{kJ}}{\text{mol}}$, $E_{4}=20\frac{\text{kJ}}{\text{mol}}$ and $E_{5}=80\frac{\text{kJ}}{\text{mol}}$.
As an example for a small network, consider the $N_{v}=5$ network depicted in Fig. 13a. This network has two spanning trees rooted at vertex $4$ (see Fig. 13b:
$Z_{\mathcal{T}}=(k_{1}+k_{2})k_{3}k_{4}k_{5},$ (39)
and four two-tree spanning forests, where one tree is rooted at vertex $4$ and the other contains vertex 0 (see Fig. 13c:
$Z_{\mathcal{F}}=k_{1}k_{4}(k_{3}+k_{5})+k_{2}k_{3}(k_{4}+k_{5}).$ (40)
Thus, the MFPT from 0 to $4$ reads:
$\langle\tau\rangle=\frac{Z_{\mathcal{F}}}{Z_{\mathcal{T}}}=\frac{1}{k_{1}+k_{2}}\Big{(}\frac{k_{1}}{k_{3}}+\frac{k_{2}}{k_{4}}\Big{)}+\frac{1}{k_{5}},$ (41)
which can give rise to the dependence in the Arrhenius plot shown in Fig. 13d for realistic values for the rate constants and activation energies.
The resulting temperature response no longer resembles the linear slope of the Arrhenius equation; instead, it is strongly curved, reaching a minimum around $T=35^{\circ}C$ and a maximum around $T=18^{\circ}C$. This implies that the global rate would actually decrease with increasing temperature within the temperature range typically relevant to biological systems. This example illustrates that the temperature dependence in medium-sized networks depends on microscopic details and may generally be more complex.
An interesting case are networks with a single rate-limiting step, i.e. a single reaction that proceeds significantly slower than the rest of the reaction network. In this case, one might naively expect that the Arrhenius equation for this reaction dominates and that the global response becomes linear again. However, this is not necessarily the case. Consider the minimal example shown in Fig. 14.
Figure 14: A minimal example to show that a network with a rate-limiting step does not necessarily yield a global Arrhenius equation dominated by this step.
Observe that:
$\langle\tau\rangle=\frac{Z_{\mathcal{F}}}{Z_{\mathcal{T}}}=\frac{k+r+\gamma}{k\gamma}=(1+\frac{r}{k})\frac{1}{\gamma}+\frac{1}{k}.$ (42)
If the step from $1$ to $2$ is rate-limiting, i.e., $\gamma\ll k,r$, then:
$\langle\tau\rangle\approx(1+\frac{r}{k})\frac{1}{\gamma},$ (43)
not just $\frac{1}{\gamma}$ as one could naively expect. So, the ratio $\frac{r}{k}\sim e^{\beta\Delta E}$ also contributes even if $\gamma$ is many orders of magnitude below $k$ and $r$. The reason for this is that temperature also impacts the equilibrium distribution. If $\gamma$ is the rate-limiting step, one can assume that the states 0 and $1$ are effectively in their equilibrium configuration, meaning that the flow from 0 to $1$ must be equal to the flow from $1$ to 0. Denoting the probability to be in state $1$ as $p$, this means:
$k(1-p)=rp\Rightarrow p=\frac{k}{k+r}.$ (44)
The crucial point is that the system can only jump to $2$ via $1$, and the rate must therefore be multiplied by the probability $p$ to be in $1$. Hence, one gets:
$\langle\tau\rangle=(p\gamma)^{-1}=(1+\frac{r}{k})\frac{1}{\gamma},$ (45)
which is the same result as before. One can therefore conclude that separation of time scales is not sufficient to obtain a pure Arrhenius dependence.
II.8.2 Linear chain
The generic quadratic behaviour predicted here can be violated not only for small networks, but also for large networks which contain only few spanning trees. The simplest possible example is a linear chain of irreversible reactions, i.e., $k_{i,j}=k_{i}\delta_{i+1,j}$, as illustrated in Fig. (15)
$\sum_{\mathcal{T}_{[N+1]}}w(\mathcal{T})=\prod_{i=1}^{N}k_{i}.$ (46)
Figure 16: Spanning forests of the linear chain. For $i\leq j$, the only corresponding spanning forest is the one without the $k_{j}$-edge. For $i>j$, there is no such tree.
Then,
One finds for the first two moments:
$\langle\tau_{i}\rangle$ $=\sum_{j=i}^{N}\frac{1}{k_{j}},$ (49)
which yields the standard deviation:
$\sigma_{\tau_{1}}=\Big{(}\sum_{j=1}^{N}\frac{1}{k_{j}^{2}}\Big{)}^{\frac{1}{2}}$ (51)
and the coefficient of variation (CV) reads:
If $k_{i}=k$ $\forall i$, one finds in general:
$\langle\tau^{n}\rangle_{i}=\frac{1}{k^{n}}\frac{(N+n-i)!}{(N-i)!},$ (53)
which are precisely the moments for an Ni+1N-i+1-dimensional Erlang distribution ($\Gamma$-distribution with non-negative integer as shape parameter) with $\lambda=\frac{1}{k}$:
$f^{\tau_{i}}(t)=\frac{k^{N-i+1}}{(N-i+1)!}t^{N-i}e^{-kt},$ (54)
as it should be since this is precisely the distribution that arises from summing up NiN-i iid exponentially distributed random variables.
Note that the coefficient of variation (CV) reads:
Moreover, the mean first-passage time for the whole process, $\langle\tau\rangle=\langle\tau_{0}\rangle$, given in Eq. (49) may be written in the limit of large $N$ as:
$\langle\tau\rangle$ $=N\langle k_{j}^{-1}\rangle=N\langle{k^{0}_{j}}^{-1}e^{\Delta\beta E_{j}}\rangle,$ (56)
writing $k_{j}$ as in Eq. (22). Assuming $k^{0}_{j}$ and $E_{j}$ to be independent, one finds:
$\ln{\langle\tau\rangle}$ $=\ln{\langle e^{\Delta\beta E_{j}}\rangle}+\ln{N}+\langle{k^{0}_{j}}^{-1}\rangle$ (57)
$=\ln{\langle e^{\Delta\beta E_{j}}\rangle}+const.$ (58)
Compare this to Eq. (36): for the linear chain, there was only a single spanning tree, so $\kappa_{1}^{E_{\mathcal{T}}}=E_{\mathcal{T}}$ and $\kappa_{n}^{E_{\mathcal{T}}}=0$ for $n\geq 0$. The total activation energies of the forests $\mathcal{F}_{i}$ for i=1,,Ni=1,...,N read:
$E_{\mathcal{F}_{i}}=\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}E_{j}=E_{\mathcal{T}}-E_{i}.$ (59)
One then obtains for the cumulant-generating function of $E_{\mathcal{F}}$:
$\sum\frac{\Delta\beta^{n}}{n!}\kappa^{E_{\mathcal{F}}}_{n}$ $=\ln{\langle e^{\Delta\beta E_{\mathcal{F}}}\rangle}$ (60)
$=\ln{e^{\Delta\beta E_{\mathcal{T}}}+\ln{\langle e^{-\Delta\beta E_{i}}\rangle}}$ (61)
$=\Delta\beta E_{\mathcal{T}}+\sum_{n=1}^{\infty}\frac{\Delta\beta^{n}}{n!}\kappa^{E_{i}}_{n}.$ (62)
Inserting this into Eq. (36) yields:
$\ln{\langle\tau\rangle}$ $=\Delta\beta E_{\mathcal{T}}+\sum_{n=1}^{\infty}\frac{\Delta\beta^{n}}{n!}\kappa^{E_{i}}_{n}-\Delta\beta E_{\mathcal{T}}+const.$ (63)
$=\ln{\langle e^{\Delta\beta E_{j}}\rangle}+const.,$ (64)
so indeed the same result as Eq. (58).
The total activation energy of the forests only differs from the total activation energy of the trees by a single $E_{i}$ and only the cumulant generating function of $E_{j}$ but not a sum over many activation energies remains. Therefore, there is no longer a statistical limit $N\to\infty$ for the distribution and the expression will depend on the distribution of $E_{i}$. If $E_{i}$ is itself normal distributed, the quadratic dependence in Eq. (37) is recovered. But choosing a uniform distribution $E_{i}\sim U([E_{\text{min}},E_{\text{max}}])$, one obtains:
$\langle e^{\Delta\beta E_{j}}\rangle$ $=\frac{1}{\Delta E}\int_{E_{\text{min}}}^{E_{\text{max}}}\text{d}E_{j}e^{\Delta\beta E_{j}}$ (65)
$=\frac{1}{\Delta E\Delta\beta}(e^{\Delta\beta E_{\text{max}}}-e^{\Delta\beta E_{\text{min}}}),$ (66)
with $\Delta E:=E_{\text{max}}-E_{\text{min}}$. Then Eq. (58) reads:
$\ln{\langle\tau\rangle}$ $=-\ln{\Delta E\Delta\beta}+\Delta\beta E_{\text{min}}+\ln{(e^{\Delta\beta\Delta E}-1)}$
$=\Delta\beta\frac{E_{\text{max}}+E_{\text{min}}}{2}+\sum_{n=2}^{\infty}\frac{\Delta E^{n}}{n!}\frac{B_{n}}{n}\Delta\beta^{n},$ (67)
where $B_{n}$ is the $n$-th Bernoulli number. This corresponds to the simulations performed by Crapse et al. [24], involving a chain of $1.000$ reactions with activation energies drawn from a uniform distribution. The authors found that the results of their simulation could not reproduce the quadratic shape that they found describing their experimental data and therefore discarded the possibility that it was a result of the complexity of the network. In light of the analytical theory developed here, that choice seems unfortunate because the linear chain is too simplistic to allow for the large system limit. Additionally, if they had chosen $E_{i}\sim\mathcal{N}(\langle E\rangle,\sigma^{2})$, it would have resulted in a quadratic shape.